HETAV PATEL

Grade 11,

a capacitor of capacitance 2nF is connected in the tank circuit of an oscillator oscillating with a frequency of 1kHz. if the current flowing in the circuit is 2mA, the voltage across the capacitor will be.

a capacitor of capacitance 2nF is connected in the tank circuit of an oscillator oscillating with a frequency of 1kHz. if the current flowing in the circuit is 2mA, the voltage across the capacitor will be.

Grade:12

1 Answers

Eshan
5 years ago
Dear student

The impedance of a capacitor=X_C=\dfrac{1}{\omega C}=\dfrac{1}{2\pi fC}=\dfrac{1}{2\pi \times 1000\times 2\times 10^{-9}}\Omega=\dfrac{1}{4\pi}\times 10^6\Omega

Hence voltage across the capacitor=V_C=iX_C=2\times 10^{-3}\times \dfrac{1}{4\pi}\times 10^6V=159.15V